/edugen.wi uni M Gmail ACE 10e 235 235 Units Question Attempts: 0 of 3 used
Answer #1

Energy produced by 1 atom of U-235 is 200 MeV
1 MeV = 1.6 * 10^-13 J
200 MeV = 3.2 * 10^-11 J

Number of moles 1 kg of U-235 will contain:
1000 / 235 = 4.255
Number of atoms = 4.255 * 6.022 * 10^23 = 2.562 * 10^24
Number of atoms in 4.6 kg = 4.6* 2.562 * 10^24 = 1.18*10^25

Total energy released = Number of atoms * Energy released per atom

(1.18* 10^25) * (3.2 * 10^-11) = 3.8 * 10^14 J

Kg of coal required to produce this amount of energy = (3.8 * 10^14) / (3 * 10^7) = 1.27 * 10^7 kgs